# Higher Order Structures For Graph Explanations

### v5

@@ 1 Introduction
Figure 1: (a) Ground Truth for an example from BENZENE (Agarwal et al. 2023) (b) Explanation generated by GNNExplainer (Ying et al. 2019) (c) Explanation generated by using FORGE. Green nodes and edges signify the subgraph considered important for GNN prediction. By incorporating FORGE, we are able to capture important multi-node interactions, resulting in more accurate explanations.
@@
However, graph explainers and GNNs face an inherent limitation in their ability to generate representations stemming from graphs’ ability to model only pair-wise interactions between nodes. This limitation has implications on the expressivity of GNNs, which is explored in detail by Xu et al. (2019). To address this, higher-order structures such as cell complexes (Hansen and Ghrist 2019) have been employed as they are higher-order generalisation of graphs, capable of modelling multi-node interactions. Neural networks designed for these higher-order structures (Bodnar et al. 2021; Ebli, Defferrard, and Spreemann 2020; Yang, Sala, and Bogdan 2022) have demonstrated significant performance improvements in graph learning tasks compared to traditional graph structures, and have been shown to be more expressive than traditional GNNs. Despite these advancements in higher-order graph representations, there exists a critical gap in the field of graph explainability: the potential of higher-order structures to enhance the interpretability of graph-based models remains unexplored.
@@
In this work, we present Framework For Higher-Order Representations In Graph Explanations (FORGE), a novel, explainer-agnostic framework designed to enhance the capability of graph explainers by capturing multi-node interactions during the GNN learning process itself by internally representing the underlying graph data as a cell complex. To further refine the quality of explanations, as part of our framework we introduce Information Propagation algorithms that translate the explanations generated for cell complexes back to the underlying graph data, resulting in richer, more accurate, and faithful explanations with respect to the input data as well as the underlying GNN predictor. This is encapsulated by Figure 1, where our framework is able to capture the multi-node interactions of a benzene ring, which is the correct ground truth explanation. Without FORGE, the base explainer is unable to interpret these interactions, resulting in less accurate explanations.
@@
To the best of our knowledge, our work marks the first exploration of the efficacy of higher-order structures in creating more interpretable representations. The overall framework is described in Figure 2. We conduct extensive evaluations of FORGE on real-world datasets from the GraphXAI (Agarwal et al. 2023) benchmark, as well as on specially curated synthetic datasets. Our results demonstrate that incorporating FORGE consistently matches or improves the explanation accuracy and faithfulness of various graph explainers. To reaffirm our hypothesis, we perform rigorous ablations and find that each components of our framework contribute significantly to increased explainer performance. Additionally, our analysis on scalability reveals that FORGE efficiently handles dense, complex graph networks with only linear overhead in both time and space complexity, further emphasising its practical applicability and computational efficiency.
@@ 2 Related Work
Hansen and Ghrist (2019) first explored relating higher-order structures to spectral graph theory. Following this, recent advancements in GNN architecture have explored the incorporation of higher-order topological structures, such as simplicial complexes (Ebli, Defferrard, and Spreemann 2020) and cell complexes (Bodnar et al. 2021), to succesfully enhance the expressive power and representational capacity of these models. This line of work is motivated by the inherent limitations of traditional GNNs, particularly their constraints in capturing complex structural patterns (Xu et al. 2019) and effectively modeling long-range dependencies (Chen, Schulz, and Borgwardt 2024).
Figure 2: Visual representation of FORGE. The input graph is lifted to a cell complex, which is then given as input to (i) a GNN to train on, as well as to (ii) a graph explainer. Propagation is then done on the output cell complex explanation to map it to an explanation for the original graph. The green color on cells, nodes and edges signify the substructure considered important for GNN prediction (the explanation).
@@
Perturbation-based methods, such as GNNExplainer (Ying et al. 2019), PGExplainer (Luo et al. 2020), SubgraphX (Yuan et al. 2021), and GraphMask (Schlichtkrull, Cao, and Titov 2022), identify important subgraphs by systematically perturbing input graphs and analyzing the impact on model outputs. Gradient-based approaches, including Grad-CAM (Selvaraju et al. 2019) and Guided Backpropagation (Springenberg et al. 2015), leverage gradient information to assess the influence of input features on model predictions. Surrogate methods, exemplified by PGM-Explainer (Vu and Thai 2020) and GraphLIME (Huang et al. 2020), approximate complex GNN behavior with interpretable local models. Recent innovations have further expanded the landscape of GNN explainability. RG-Explainer (Shan et al. 2021) employs reinforcement learning to construct explanatory subgraphs tailored to both model architecture and individual instances. MATE (Spinelli, Scardapane, and Uncini 2024) introduces a meta-learning framework that jointly optimises GNN performance and explainability, resulting in intrinsically more interpretable representations.
While these approaches have significantly advanced GNN explainability, they primarily operate on traditional graph structures. Our proposed framework, FORGE, distinguishes itself by leveraging higher-order topological structures, specifically cell complexes, (1) to train GNNs to inherently create more interpretable representations during the learning process, and (2) utilise their topological properties to refine explanations post-hoc.
@@ 3 Cell Complexes
Definition 3.2 (boundary/coboundary).
@@
Definition 3.5 (boundary relation).
@@
While cell complexes are adept at handling higher-order interactions, their conventional definition requires them to be closed under taking subsets, making them inefficient data structures for scalable computation (Yang, Sala, and Bogdan 2022). To overcome this, we present the following theorem and a subsequent modification in Equation (3) to achieve an efficient, restricted form of cell complexes that is linear in space complexity. The proof follows by construction, and is deferred to the Appendix.
@@
This theorem provides us two important bounds. Equation (1) shows that the number of cells in cell complex $X$ grows linearly with the number of vertices and edges in the graph $G$. Equation (2) shows that asymptotically, the size of $\Sigma$ grows quadratically with $|E|$, due to the last term in the equation. This term is precisely the number of boundary relations that are present between two edges in the graph, which are represented by $\sigma_{c_{1}^{(1)},c_{2}^{(1)}}$. To preserve the scalabilty and complexity of FORGE, we further restrict how we create cell complexes from graphs, and drop all boundary relations of the form $\sigma_{c_{1}^{(p)},c_{2}^{(p)}}$ where $p\geq 1$ from our construction of cell complexes. Doing this reduces Equation (2) to:
@@ 4 Proposed Approach
In this section, we present the lifting algorithm we introduce to create a cell complex $X$ from its underlying graph $G$. Using the lifting algorithm on a given graph $G(V,E)$, we construct $C^{(0)}$, $C^{(1)}$, and $C^{(2)}$ from $V$, $E$, and cycles present in the graph respectively, and add the corresponding $\Sigma$ to create $X(C,\Sigma)$. The lifting algorithm is described in Algorithm 1. For any given graph, the time complexity of the lifting algorithm is bounded by the time complexity of finding cycles in a graph [^2] https://networkx.org/documentation/stable/reference/algorithms/generated/networkx.algorithms.cycles.simple˙cycles.html. Further details on the algorithm and how cell features are created in order to perform message passing are deferred to the Appendix.
@@
While graph explainers use diverse methods to generate explanations for GNN predictions, they can be abstracted as a function that takes in the trained GNN model and graph data, and outputs a soft node or edge mask with values between 0 and 1 signifying the importance of each node and edge identified as the explanation. For simplicity, we continue the discussion by defining the explanation as a node mask, an analogous approach can be followed for an edge mask. Formally, let $\mathcal{M}$ be an explanation mask, then the output of a graph explainer $\mathcal{E}$ can be represented as
@@
Graph explainers generate explanations on the computation graph of a node, which is the node’s $k$-hop neighbourhood for a $k$-layer GNN. Transitioning from traditional computation graphs, our approach utilises cell complexes as the input to the function. This shift allows the integration of more complex structural data into the computational framework, through the introduction of vertical message passing happening through vertical boundary relations (Section 3) in the cell complex, in addition to the already existing (horizontal) message passing, shown in Figure 4. Consequently, the learned explanation mask is adapted to represent the computational complex, incorporating both the connectivity and hierarchical structure inherent in cell complexes. Formally, this can be represented by
@@
We introduce multiple algorithms to perform information propagation, namely, (1) Hierarchical Propagation, (2) Direct Propagation, (3) Entropy Propagation and (4) Activation Propagation. For brevity, we expand on (1) in detail, and defer the description of other algorithms and optimisations to the Appendix. Hierarchical Propagation is described in Algorithm 2.
@@
The main idea behind Hierarchical Propagation is to aggregate the explanations of 2-cells with their 1-cell boundaries, then again aggregate explanations of 1-cells with their 0-cell boundaries. These 0-cells represent underlying 00-skeleton of $X$, which as mentioned in Section 3 is precisely the set of nodes $V$ of $G(V,E)$. The goal is to utilise cell explanations in a way that polarises important and unimportant nodes in the final explanation mask $\mathcal{M}$. This is done in Algorithm 2 in steps 3 and 8, where we subtract $0.5$ from the explanation masks of higher order cells in order to quantify how positively or negatively it should influence the overall explanation. They are further multiplied by parameters $\alpha_{e}$ and $\alpha_{c}$, which are introduced to have fine-grained control over how much 1-cells and 2-cells, respectively, contribute to the graph explanation. By precomputing the required boundary relations, the time complexity of Information Propagation is reduced to $\mathcal{O}(|E|)$.
@@ 5 Experimental Settings
Table 2: Graph Explanation Accuracy (GEA) ($\uparrow$) scores for baseline explainers (B) against their FORGE variants (F) across all datasets, with FORGE improving performance across various baselines. Best result for each dataset is highlighted in bold. Underlined values represent the better result between a base explainer and its FORGE variant.
@@
We evaluate FORGE on both real-world and synthetic datasets for a comprehensive evaluation across diverse conditions. We take real-world datasets from the graph explainability benchmark, GraphXAI (Agarwal et al. 2023), which includes Benzene, Mutagenicity, Alkyl Carbonyl, and Fluoride Carbonyl. For synthetic datasets, we generate random graphs with a distinct subgraph structure, known as a motif, which defines the ground truth for the graph. The task involves differentiating between two different motifs. Motifs we test on include Bull, Square (4-Cycle), Hexagon (6-Cycle), Wheel, House, and Cube (Figure 5. The synthetic datasets are referred to as Motif1/Motif2 in the text based on the motifs present. Specific details about dataset generation can be found in the Appendix.
Figure 5: Different motifs used to generate synthetic graphs: (left to right) Bull, Wheel, Cube, and House
@@ 6 Results
Table 2 presents the GEA scores for various explainers with and without our framework. The reported results are averaged over $10$ different seeds, with standard deviations reported in the Appendix. Our experiments reveal several key insights into the performance of different GNN explainers across various datasets. FORGE consistently improves existing explainers across various datasets with improvements up to $315$% (FORGE-enhanced GradExplainer on Cube/Wheel), with the exception of Mutag for GNNExplainer. The GEF scores are presented in Table 3, reinforcing the capabilities of our framework in creating explanations that are comparable or more faithful to the underlying GNN.
@@
Table 3: Graph Explanation Faithfulness (GEF) Scores for the AlkaneCarbonyl Dataset. FORGE enhanced explainers generate explanations that are comparable or more faithful to the underlying GNN. Values in bold indicate best performance.
@@ 7 Ablation Studies
FORGE comprises two key components: Lifting and Information Propagation. To evaluate their individual contributions, we compare FORGE’s performance against FORGE-LIFT (a version that only performs Lifting) and the baseline explainer (which has neither Lifting nor Propagation). Figure 5 illustrates the results across four different graph explainers. Figure 6 demonstrate that FORGE-LIFT consistently outperforms the base explainers across all four methods. This improvement underscores the significant impact of the Lifting component alone in generating more accurate explanations. Furthermore, when we incorporate Information Propagation to create the full FORGE framework, we observe an additional performance boost on top of Lifting. This enhancement is particularly noticeable in GNNExplainer and GradExplainer (subfigures (a) and (c)).
Figure 6: Results of ablations on FORGE Components for (a) GNNExplainer, (b) GraphMaskExplainer, (c) GradExplainer, (d) PGMExplainer on Synth-Wheel/Cube dataset. Both Lifting and Information Propagation contribute significantly to increase in explanation accuracy.
@@
Figure 7 presents our second ablation study, comparing four information propagation algorithms across four different datasets for all explainers, reporting the average GEA. This analysis reveals two key insights: (a) Algorithm Performance Variability: the effectiveness of propagation methods varies significantly across datasets, indicating that no single algorithm consistently outperforms the others in all scenarios, and (b) Dataset Dependency: the optimal choice of propagation algorithm seems to be heavily influenced by the specific characteristics of the dataset being analysed. This suggests that the graph structure and data properties play an important role in determining the most effective propagation method.

### v6

@@ 1 Introduction
Figure 1: (a) Ground Truth for an example from BENZENE (Agarwal et al. 2023) (b) Explanation generated by GNNExplainer (Ying et al. 2019) (c) Explanation generated by using FORGE. Green nodes and edges signify the subgraph considered important for GNN prediction. By incorporating FORGE, we can capture important multi-node interactions, resulting in more accurate explanations.
@@
However, graph explainers and GNNs face an inherent limitation in their ability to generate representations stemming from graphs’ ability to model only pair-wise interactions between nodes. This limitation has implications on the expressivity of GNNs, which is explored in detail by Xu et al. (2019). To address this, higher-order structures such as cell complexes (Hansen and Ghrist 2019) have been employed as they are a higher-order generalization of graphs, capable of modeling multi-node interactions. Neural networks designed for these higher-order structures (Bodnar et al. 2021; Ebli, Defferrard, and Spreemann 2020; Yang, Sala, and Bogdan 2022) have demonstrated significant performance improvements in graph learning tasks compared to traditional graph structures, and have been shown to be more expressive than traditional GNNs. Despite these advancements in higher-order graph representations, there exists a critical gap in the field of graph explainability: the potential of higher-order structures to enhance the interpretability of graph-based models remains unexplored.
@@
In this work, we present Framework For Higher-Order Representations In Graph Explanations (FORGE), a novel, explainer-agnostic framework designed to enhance the capability of graph explainers by capturing multi-node interactions during the GNN learning process itself by internally representing the underlying graph data as a cell complex. To further refine the quality of explanations, as part of our framework we introduce Information Propagation algorithms that translate the explanations generated for cell complexes back to the underlying graph data, resulting in richer, more accurate, and faithful explanations with respect to the input data as well as the underlying GNN predictor. This is encapsulated by Figure 1, where our framework can capture the multi-node interactions of a benzene ring, which is the correct ground truth explanation. Without FORGE, the base explainer is unable to interpret these interactions, resulting in less accurate explanations.
@@
The overall framework is described in Figure 2. We conduct extensive evaluations of FORGE on real-world datasets from the GraphXAI (Agarwal et al. 2023) benchmark, as well as on specially curated synthetic datasets. Our results demonstrate that incorporating FORGE consistently matches or improves the explanation accuracy and faithfulness of various graph explainers. To reaffirm our hypothesis, we perform rigorous ablations and find that each components of our framework contribute significantly to increased explainer performance. Additionally, our analysis on scalability reveals that FORGE efficiently handles dense, complex graph networks with only linear overhead in both time and space complexity, further emphasizing its practical applicability and computational efficiency.
@@ 2 Related Work
Hansen and Ghrist (2019) first explored relating higher-order structures to spectral graph theory. Following this, recent advancements in GNN architecture have explored the incorporation of higher-order topological structures, such as simplicial complexes (Ebli, Defferrard, and Spreemann 2020) and cell complexes (Bodnar et al. 2021), to successfully enhance the expressive power and representational capacity of these models. This line of work is motivated by the inherent limitations of traditional GNNs, particularly their constraints in capturing complex structural patterns (Xu et al. 2019) and effectively modeling long-range dependencies (Chen, Schulz, and Borgwardt 2024).
Figure 2: Visual representation of FORGE. The input graph is lifted to a cell complex, which is then given as input to (i) a GNN to train on, as well as to (ii) a graph explainer. Propagation is then done on the output cell complex explanation to map it to an explanation for the original graph. The green color on cells, nodes, and edges signify the substructure considered important for GNN prediction (the explanation).
@@
Perturbation-based methods, such as GNNExplainer (Ying et al. 2019), PGExplainer (Luo et al. 2020), SubgraphX (Yuan et al. 2021), and GraphMask (Schlichtkrull, Cao, and Titov 2022), identify important subgraphs by systematically perturbing input graphs and analyzing the impact on model outputs. Gradient-based approaches, including Grad-CAM (Selvaraju et al. 2019) and Guided Backpropagation (Springenberg et al. 2015), leverage gradient information to assess the influence of input features on model predictions. Surrogate methods, exemplified by PGM-Explainer (Vu and Thai 2020) and GraphLIME (Huang et al. 2020), approximate complex GNN behavior with interpretable local models. Recent innovations have further expanded the landscape of GNN explainability. RG-Explainer (Shan et al. 2021) employs reinforcement learning to construct explanatory subgraphs tailored to both model architecture and individual instances. MATE (Spinelli, Scardapane, and Uncini 2024) introduces a meta-learning framework that jointly optimizes GNN performance and explainability, resulting in intrinsically more interpretable representations. MotifExplainer (Yu and Gao 2022) introduces a motif-based approach for human interpretable explanations. While MotifExplainer is most similar to our work, it is domain specific and suffers from $O(n^{3})$ complexity.
Although these approaches have significantly advanced GNN explainability, they primarily operate on traditional graph structures. Our proposed framework, FORGE, distinguishes itself by leveraging higher-order topological structures, specifically cell complexes, (1) to train GNNs to inherently create more interpretable representations during the learning process, and (2) utilise their topological properties to refine explanations post-hoc.
@@ 3 Cell Complexes
Definition 2 (boundary/coboundary).
@@
Definition 5 (boundary relation).
@@
While cell complexes are adept at handling higher-order interactions, their conventional definition requires them to be closed under taking subsets, making them inefficient data structures for scalable computation (Yang, Sala, and Bogdan 2022). To overcome this, we present the following theorem and a subsequent modification in Equation (3) to achieve an efficient, restricted form of cell complexes that is linear in space complexity. The proof follows by construction and is deferred to the Appendix.
@@
This theorem provides us with two important bounds. Equation (1) shows that the number of cells in cell complex $X$ grows linearly with the number of vertices and edges in the graph $G$. Equation (2) shows that asymptotically, the size of $\Sigma$ grows quadratically with $|E|$, due to the last term in the equation. This term is precisely the number of boundary relations that are present between two edges in the graph, which are represented by $\sigma_{c_{1}^{(1)},c_{2}^{(1)}}$. To preserve the scalability and complexity of FORGE, we further restrict how we create cell complexes from graphs, and drop all boundary relations of the form $\sigma_{c_{1}^{(p)},c_{2}^{(p)}}$ where $p\geq 1$ from our construction of cell complexes. Doing this reduces Equation (2) to:
@@ 4 Proposed Approach
In this section, we present the lifting algorithm we introduce to create a cell complex $X$ from its underlying graph $G$. Using the lifting algorithm on a given graph $G(V,E)$, we construct $C^{(0)}$, $C^{(1)}$, and $C^{(2)}$ from $V$, $E$, and cycles present in the graph respectively, and add the corresponding $\Sigma$ to create $X(C,\Sigma)$. The lifting algorithm is described in Algorithm 1. For any given graph, the time complexity of the lifting algorithm is bounded by the time complexity of finding cycles in a graph [^2] https://networkx.org/documentation/stable/reference/algorithms/generated/networkx.algorithms.cycles.simple˙cycles. Further details on the algorithm and how cell features are created to perform message passing are deferred to the Appendix.
@@
While graph explainers use diverse methods to generate explanations for GNN predictions, they can be abstracted as a function that takes in the trained GNN model and graph data and outputs a soft node or edge mask with values between 0 and 1 signifying the importance of each node and edge identified as the explanation. For simplicity, we continue the discussion by defining the explanation as a node mask, an analogous approach can be followed for an edge mask. Formally, let $\mathcal{M}$ be an explanation mask, then the output of a graph explainer $\mathcal{E}$ can be represented as
@@
Graph explainers generate explanations on the computation graph of a node, which is the node’s $k$-hop neighborhood for a $k$-layer GNN. Transitioning from traditional computation graphs, our approach utilizes cell complexes as the input to the function. This shift allows the integration of more complex structural data into the computational framework, through the introduction of vertical message passing happening through vertical boundary relations (Section 3) in the cell complex, in addition to the already existing (horizontal) message passing, shown in Figure 4. Consequently, the learned explanation mask is adapted to represent the computational complex, incorporating both the connectivity and hierarchical structure inherent in cell complexes. Formally, this can be represented by
@@
We introduce multiple algorithms to perform information propagation, namely, (1) Hierarchical Propagation, (2) Direct Propagation, (3) Entropy Propagation, and (4) Activation Propagation. For brevity, we expand on (1) in detail and defer the description of other algorithms and optimizations to the Appendix. Hierarchical Propagation is described in Algorithm 2.
@@
The main idea behind Hierarchical Propagation is to aggregate the explanations of 2-cells with their 1-cell boundaries, then again aggregate explanations of 1-cells with their 0-cell boundaries. These 0-cells represent the underlying 00-skeleton of $X$, which as mentioned in Section 3 is precisely the set of nodes $V$ of $G(V,E)$. The goal is to utilize cell explanations in a way that polarises important and unimportant nodes in the final explanation mask $\mathcal{M}$. This is done in Algorithm 2 in steps 3 and 8, where we subtract $0.5$ from the explanation masks of higher order cells in order to quantify how positively or negatively it should influence the overall explanation. They are further multiplied by parameters $\alpha_{e}$ and $\alpha_{c}$, which are introduced to have fine-grained control over how much 1-cells and 2-cells, respectively, contribute to the graph explanation. By precomputing the required boundary relations, the time complexity of Information Propagation is reduced to $\mathcal{O}(|E|)$.
@@ 5 Experimental Settings
Table 2: Graph Explanation Accuracy (GEA) ($\uparrow$) scores for baseline explainers (B) against their FORGE variants (F) across all datasets, with FORGE improving performance across various baselines. The best result for each dataset is highlighted in bold. Underlined values represent the better result between a base explainer and its FORGE variant.
@@
We evaluate FORGE on both real-world and synthetic datasets for a comprehensive evaluation across diverse conditions. We take real-world datasets from the graph explainability benchmark, GraphXAI (Agarwal et al. 2023), which includes Benzene, Mutagenicity, Alkyl Carbonyl, and Fluoride Carbonyl. For synthetic datasets, we generate random graphs with a distinct subgraph structure, known as a motif, which defines the ground truth for the graph. The task involves differentiating between two different motifs. Motifs we test on include Bull, Square (4-Cycle), Hexagon (6-Cycle), Wheel, House, and Cube (Figure 5). The synthetic datasets are referred to as Motif1/Motif2 in the text based on the motifs present. Specific details about dataset generation can be found in the Appendix.
Figure 5: Different motifs used to generate synthetic graphs: (left to right) Bull, Wheel, Cube, and House.
@@ 6 Results
Table 2 presents the GEA scores for various explainers with and without our framework. The reported results are averaged over $10$ different seeds, with standard deviations reported in the Appendix. Our experiments reveal several key insights into the performance of different GNN explainers across various datasets. FORGE consistently improves existing explainers across various datasets with improvements up to $315$% (FORGE-enhanced GradExplainer on Cube/Wheel), except Mutag for GNNExplainer. The GEF scores are presented in Table 3, reinforcing the capabilities of our framework in creating explanations that are comparable or more faithful to the underlying GNN.
@@
Table 3: Graph Explanation Faithfulness (GEF) Scores for the AlkaneCarbonyl Dataset. FORGE-enhanced explainers generate explanations that are comparable or more faithful to the underlying GNN. Values in bold indicate best performance.
@@ 7 Ablation Studies
FORGE comprises two key components: Lifting and Information Propagation. To evaluate their individual contributions, we compare FORGE’s performance against Base-LIFT (a version that applies regular explainers on lifted graphs), FORGE-LIFT (a version of FORGE that only performs lifting), and Base (which has neither Lifting nor Propagation). Figure 6 demonstrates that FORGE-LIFT, by itself, consistently outperforms the base explainers across all four methods. While Base-LIFT performs better than the baseline, we see that training the underlying GNN on the lifted graphs is important to achieve further performance gains. This improvement highlights the significant impact of the Lifting component alone in generating more accurate explanations. Furthermore, when we incorporate Information Propagation to create the full FORGE framework, we observe an additional performance boost on top of Lifting. This enhancement is particularly noticeable in GNNExplainer and GradExplainer (subfigures (a) and (c)).
Figure 6: Results of ablations on FORGE Components for (a) GNNExplainer, (b) GraphMaskExplainer, (c) GradExplainer, (d) PGMExplainer on Synth-Wheel/Cube dataset. Both Lifting and Information Propagation contribute significantly to an increase in explanation accuracy.
@@
Figure 7 presents our second ablation study, comparing four information propagation algorithms across four different datasets for all explainers, reporting the average GEA. This analysis reveals two key insights: (a) Algorithm Performance Variability: the effectiveness of propagation methods varies significantly across datasets, indicating that no single algorithm consistently outperforms the others in all scenarios, and (b) Dataset Dependency: the optimal choice of propagation algorithm seems to be heavily influenced by the specific dataset being analyzed. This suggests that the graph structure and data properties play an important role in determining the most effective propagation method.